OpenStudy (turtlesforlife):

Match each function with the correct inverse. f(x)=1/5x+3 f(x)=5x+3 f(x)=(x+5)^2 f(x)=x^2+5

OpenStudy (turtlesforlife):

The options I have are a. \[f ^{-1}(x)=\sqrt{x-5}; Domain: [0,\infty)\]

OpenStudy (turtlesforlife):

\[f ^{-1}(x)=5(x-3)\]